OpenStudy (anonymous):

can someone help me solve multi-step equations and inequalites 5n+3=2(n+2)-3n

OpenStudy (anonymous):

5n +3 = 2(n+2) -3n so first lets look at the right hand side of the equation and multiple out the brackets. get it to be as simple as possible... 2(n+2)-3n ==> 2n +4 -3n so now we have 5n +3 = 2n + 4 -3n now add them all up so you only have one set of "n"s on either side... 5n+ 3 = 4 -n (because 2n - 3n = -n) now get the n on oneside and the numbers on the other. remember when you move numbers across the = sign on an equation you have to change their + or - sign. 5n + n = 4 - 3 6n = 1 so n = 1/6